Find the finite population correction factor given the following:




a. N= 300, n= 50




b.




N= 2500, n= 100




C.




N = 150, n= 10

1
Expert's answer
2022-05-31T08:49:05-0400

"FPC=\sqrt{\frac{N-n}{N-1}}."

  1. "FPC=\sqrt{\frac{300-50}{300-1}}=0.914"
  2. "FPC=\sqrt{\frac{2500-100}{2500-1}}=0.979"
  3. "FPC=\sqrt{\frac{150-10}{150-1}}=0.97"